# Learning with Local Search MCMC Layers

### v1

@@ B.3 Proposal distribution design for the DVRPTW
The second correction factor needed is $\frac{|Q(\bm{y})|}{|Q(\bm{y}^{\prime})|}$ (see Algorithm 2). We compute it by checking if each move is allowed, i.e., if there exists at least one $i\in V^{1}_{s}(\bm{y})$ such that $V^{2}_{s}(\bm{y})[i]\setminus\{i\}\neq\emptyset$. This can be determined in $\mathcal{O}(\mathcal{R}^{\omega})$ for each move.
@@
B.4 Additional experimental details and results for Section 5
@@
Following Baty et al. [4], the differentiable ML model $g_{W}$ is implemented as a sparse graph neural network. We use the same feature set (named complete feature set, and described in the Table 4 of their paper). We use the same training, validation, and testing datasets, which are created from $30$, $15$ and $25$ problem instances respectively. The training set uses a sample size of $50$ requests per wave, while the rest use $100$. The solutions in the training dataset, i.e., the examples from the anticipative strategy $f^{\star}$ imitated by the model, are obtained by solving the corresponding offline VRPTWs using HGS [51] with a time limit of $3600$ seconds. During evaluation, the PC-HGS solver $\tilde{{\bm{y}}}$ is used with a constant time limit of $60$ seconds for all models. We use Adam [25] together with the proposed stochastic gradient estimators, with a learning rate of $5\cdot 10^{-3}$. Each training is performed using only a single CPU worker. For Fig. 2, we use a temperature $t=10^{2}$. For Table 3, we use $1$ Monte-Carlo sample for the perturbation-based method and $1$ Markov chain for the proposed approach (in order to have a fair comparison: an equal number of oracle calls / equal compute).
@@
Each training is performed $50$ times with the same parameters and different random seeds. Then, the learning curves are averaged, and plotted with a $95\%$ confidence interval. For the results in Table 3, we report the performance of the best model iteration (selected with respect to the validation set) on the test set. This procedure is also averaged over $50$ trainings, and reported with $95\%$ confidence intervals.
@@
In Fig. 13, we report model performance for varying temperature $t$. Interestingly, lower temperatures perform better when using random initialization. In the ground-truth initialization setting, a sweet spot is found at $t=10^{2}$, but lower temperatures do not particularly decrease performance.
@@
Appendix C Proofs
C.1 Proof of Eq. 4
@@ Appendix C Proofs
At fixed temperature $t_{k}=t$, the iterates of Algorithm 1 (MH case) follow a time-homogenous Markov chain, defined by the following transition kernel $P_{\bm{\theta},t}$:
@@
$P_{\bm{\theta},t}(\bm{y},\bm{y})$ $=1-\sum_{\bm{y}^{\prime}\in\mathcal{N}(\bm{y})}P_{\bm{\theta},t}(\bm{y},\bm{y}^{\prime})$
@@
Thus, we have $P_{\bm{\theta},t}(\bm{y},\bm{y})>0$, which implies that the chain is aperiodic. As an irreducible and aperiodic Markov Chain on a finite state space, it converges to its stationary distribution and the latter is unique [18]. Finally, one can easily check that the detailed balance equation is satisfied for ${\pi}_{\bm{\theta},t}$, i.e. :
@@
$t\rightarrow+\infty]{}\frac{1}{|\mathcal{Y}|},$
as $\exp(x/t)\xrightarrow[t\rightarrow+\infty]{}1$ for all $x\in\mathbb{R}$. Thus, ${\pi}_{\bm{\theta},t}$ converges to the uniform distribution on $\mathcal{Y}$, and its expectation converges to the average of all structures.
Expression of the Jacobian. Let $A_{t}:\bm{\theta}\mapsto t\cdot\log\sum_{\bm{y}\in\mathcal{Y}}\exp\left(\langle\bm{\theta},\bm{y}\rangle+\varphi(\bm{y})\right)$ be the cumulant function of the exponential family defined by ${\pi}_{\bm{\theta},t}$, scaled by $t$. One can easily check that we have $\nabla_{\bm{\theta}}A_{t}(\bm{\theta})={\widehat{\bm{y}}_{t}}(\bm{\theta})$. Thus, we have $J_{\bm{\theta}}{\widehat{\bm{y}}_{t}}(\bm{\theta})=\nabla^{2}_{\bm{\theta}}A_{t}(\bm{\theta})$. However, we also have that the hessian matrix of the cumulant function $\bm{\theta}\mapsto\frac{1}{t}A_{t}(\bm{\theta})$ is equal to the covariance matrix of the random vector $\frac{Y}{t}$ under ${\pi}_{\bm{\theta},t}$ [53]. Thus, we have:
@@
Let $K_{\bm{\theta},t}$ be the Markov transition kernel associated to Algorithm 2, which can be written as:
$K_{\bm{\theta},t}(\bm{y},\bm{y}^{\prime})=\begin{cases}\sum_{\begin{subarray}{c}s\in Q(\bm{y})\\ \text{s.t. }q_{s}(\bm{y},\bm{y}^{\prime})>0\end{subarray}}\frac{1}{|Q(\bm{y})|}q_{s}(\bm{y},\bm{y}^{\prime})\min\left(1,\frac{|Q(\bm{y})|}{|Q(\bm{y}^{\prime})|}\cdot\frac{q_{s}(\bm{y^{\prime}},\bm{y}){\pi}_{\theta,t}(\bm{y}^{\prime})}{q_{s}(\bm{y},\bm{y}^{\prime}){\pi}_{\theta,t}(\bm{y})}\right)&\text{if $\bm{y}^{\prime}\in\bar{\mathcal{N}}(\bm{y})$,}\\ 1-\sum_{\bm{y}^{\prime\prime}\in\bar{\mathcal{N}}(\bm{y})}K_{\bm{\theta},t}(\bm{y},\bm{y}^{\prime\prime})&\text{if $\bm{y}^{\prime}=\bm{y}$,}\\ 0&\text{else.}\end{cases}$
@@
Thus, we only have to check that the detailed balance equation ${\pi}_{\theta,t}(\bm{y})K_{\bm{\theta},t}(\bm{y},\bm{y}^{\prime})={\pi}_{\theta,t}(\bm{y}^{\prime})K_{\bm{\theta},t}(\bm{y}^{\prime},\bm{y})$ is satisfied for all $\bm{y}^{\prime}\in\bar{\mathcal{N}}(\bm{y})$. We have:
${\pi}_{\theta,t}(\bm{y})K_{\bm{\theta},t}(\bm{y},\bm{y}^{\prime})=\sum_{\begin{subarray}{c}s\in Q(\bm{y})\\ \text{s.t. }q_{s}(\bm{y},\bm{y}^{\prime})>0\end{subarray}}\biggl{[}\frac{q_{s}(\bm{y},\bm{y}^{\prime}){\pi}_{\theta,t}(\bm{y})}{|Q(\bm{y})|}\min\left(1,\frac{|Q(\bm{y})|}{|Q(\bm{y}^{\prime})|}\cdot\frac{q_{s}(\bm{y^{\prime}},\bm{y}){\pi}_{\theta,t}(\bm{y}^{\prime})}{q_{s}(\bm{y},\bm{y}^{\prime}){\pi}_{\theta,t}(\bm{y})}\right)\biggr{]}.$
The main point consists in noticing that the undirectedness assumption for each neighborhood graph $G_{\mathcal{N}_{s}}$ implies:
@@
Thus, a simple case analysis on how $|Q(\bm{y})|q_{s}(\bm{y}^{\prime},\bm{y}){\pi}_{\theta,t}(\bm{y}^{\prime})$ and $|Q(\bm{y}^{\prime})|q_{s}(\bm{y},\bm{y}^{\prime}){\pi}_{\theta,t}(\bm{y})$ compare allows us to observe that the expression of ${\pi}_{\theta,t}(\bm{y})K_{\bm{\theta},t}(\bm{y},\bm{y}^{\prime})$ is symmetric in $\bm{y}$ and $\bm{y}^{\prime}$, which concludes the proof. ∎
C.4 Proof of strict convexity
@@
Moreover, as $\nabla A_{t}(\mathbb{R}^{d})=\operatorname{\mathsf{relint}}(\mathcal{C})$, we have $||\nabla A_{t}(\bm{\theta})||\leq R_{\mathcal{C}}\coloneqq\max_{\bm{\mu}\in\mathcal{C}}||\bm{\mu}||$, which gives $\operatorname{\mathsf{dom}}(\Omega_{t})\subset B(\bm{0},R_{\mathcal{C}})$. Thus we can actually write:
@@
The proposed Fenchel-Young loss can also be obtained via distribution-space regularization. Let $s_{\bm{\theta}}\coloneqq\left(\langle\bm{\theta}\,,\,\bm{y}\rangle+\varphi(\bm{y})\right)_{\bm{y}\in\mathcal{Y}}\in\mathbb{R}^{|\mathcal{Y}|}$ be a vector containing the score of all structures, and $L_{-tH}:\mathbb{R}^{|\mathcal{Y}|}\times\Delta^{|\mathcal{Y}|}\rightarrow\mathbb{R}$ be the Fenchel-Young loss generated by $-tH$, where $H$ is the Shannon entropy. We have $\nabla_{s_{\bm{\theta}}}(-tH)^{*}(s_{\bm{\theta}})={\pi}_{\bm{\theta},t}$. The chain rule further gives $\nabla_{\bm{\theta}}(-tH)^{*}(s_{\bm{\theta}})=\mathbb{E}_{{\pi}_{\bm{\theta},t}}\left[Y\right]$. Thus, we have $\nabla_{\bm{\theta}}L_{-tH}(s_{\bm{\theta}}\,;\bm{p}_{\bm{y}})=\nabla_{\bm{\theta}}\ell_{t}(\bm{\theta}\,;\bm{y})$, where $\bm{p}_{\bm{y}}$ is the dirac distribution on $\bm{y}$. In the case where $\varphi\equiv 0$ and $t=1$, we have $\Omega_{t}(\bm{\mu})=-\left(\max_{\bm{p}\in\Delta^{|\mathcal{Y}|}}H^{s}(\bm{p})\text{ s.t. }\mathbb{E}_{\bm{p}}\left[Y\right]=\bm{\mu}\right)$, with $H^{s}$ the Shannon entropy [8], and $\ell_{t}$ is known as the CRF loss [29].
C.5 Proof of Section 4.4
@@
Moreover, we already derived $\nabla^{2}_{\bm{\theta}}A_{t}(\bm{\theta}_{0})=\frac{1}{t}\operatorname{\mathsf{cov}}_{{\pi}_{\bm{\theta}_{0},t}}\left[Y\right]$ in Section C.2, leading to the simplified asymptotic normality given in the proposition.
C.6 Proof of Section 4.4
@@
More precisely, we must find a constant $D$ such that $\exists B>0\,:\;\lambda_{\bm{\theta}}\leq 1-Be^{-D||\bm{\theta}||}$, in order to impose $K_{n+1}>\left\lfloor 1+a^{\prime}\exp\left(2D||\hat{\bm{\theta}}_{n}||\right)\right\rfloor$.
A known result gives $\lambda_{\bm{\theta}}\leq\rho(\bm{\theta})$ with $\rho(\bm{\theta})=\max_{\lambda\in S_{\bm{\theta}}\backslash\{1\}}|\lambda|$ [32], where $S_{\bm{\theta}}$ is the spectrum of the transition kernel $P_{\bm{\theta},t}$ (here, $1-\rho(\bm{\theta})$ is known as the spectral gap of the Markov chain). To bound $\rho(\bm{\theta})$, we use the results of Ingrassia [23], which study the Markov chain with transition kernel $P^{\prime\varphi}_{\bm{\theta},t}$, such that $P_{\bm{\theta},t}=\frac{1}{2}\left(I+P^{\prime\varphi}_{\bm{\theta},t}\right)$. It corresponds to the same algorithm, but with a proposal distribution $q^{\prime}$ defined as:
@@
As $P^{\prime\varphi}_{\bm{\theta},t}$ is a row-stochastic matrix, Gershgorin’s circle theorem gives that its spectrum is included in the complex unit disc. Moreover, as one can easily check that the associated Markov chain also has ${\pi}_{\bm{\theta},t}$ as stationary distribution, the detailed balance equation gives:
$\forall\bm{y},\bm{y}^{\prime}\in\mathcal{Y},\;{\pi}_{\bm{\theta},t}(\bm{y})P^{\prime\varphi}_{\bm{\theta},t}(\bm{y},\bm{y}^{\prime})={\pi}_{\bm{\theta},t}(\bm{y}^{\prime})P^{\prime\varphi}_{\bm{\theta},t}(\bm{y}^{\prime},\bm{y}),$
@@
$\forall\bm{y},\bm{y}^{\prime}\in\mathcal{Y},\;\sqrt{\frac{{\pi}_{\bm{\theta},t}(\bm{y})}{{\pi}_{\bm{\theta},t}(\bm{y}^{\prime})}}P^{\prime\varphi}_{\bm{\theta},t}(\bm{y},\bm{y}^{\prime})=\sqrt{\frac{{\pi}_{\bm{\theta},t}(\bm{y}^{\prime})}{{\pi}_{\bm{\theta},t}(\bm{y})}}P^{\prime\varphi}_{\bm{\theta},t}(\bm{y}^{\prime},\bm{y})$
@@
$\Pi_{\bm{\theta}}^{1/2}P^{\prime\varphi}_{\bm{\theta},t}\Pi_{\bm{\theta}}^{-1/2}=\Pi_{\bm{\theta}}^{-1/2}P^{\prime\varphi\top}_{\bm{\theta},t}\Pi_{\bm{\theta}}^{1/2},$
where $\Pi_{\bm{\theta}}=\operatorname{\mathsf{diag}}({\pi}_{\bm{\theta};t})$. Thus, the matrix $\Pi_{\bm{\theta}}^{1/2}P^{\prime\varphi}_{\bm{\theta},t}\Pi_{\bm{\theta}}^{-1/2}$ is symmetric, and the spectral theorem ensures its eigenvalues are real. As it is similar to the transition kernel $P^{\prime\varphi}_{\bm{\theta},t}$ (with change of basis matrix $\Pi_{\bm{\theta}}^{-1/2}$), they share the same spectrum $S^{\prime}_{\bm{\theta}}$, and we have $S^{\prime}_{\bm{\theta}}\subset\left[-1,1\right]$. Let us order $S^{\prime}_{\bm{\theta}}$ as $-1\leq\lambda^{\prime}_{\min}\leq\dots\leq\lambda^{\prime}_{2}\leq\lambda^{\prime}_{1}=1$. As $P_{\bm{\theta},t}=\frac{1}{2}\left(I+P^{\prime\varphi}_{\bm{\theta},t}\right)$, we clearly have $\rho(\bm{\theta})=\frac{1+\lambda^{\prime}_{2}}{2}$. Thus, we can use Theorem 4.1 of Ingrassia [23], which gives $\lambda^{\prime}_{2}\leq 1-G\cdot Z(\bm{\theta})\exp(-m\left(\bm{\theta})\right)$ (we keep their notations for $Z$ and $m$, and add the dependency in $\bm{\theta}$ for clarity), where $G$ is a constant depending only on the graph $G_{\mathcal{N}}$, and with:
@@
$\geq|\mathcal{Y}|\exp\left(-\frac{2R_{\mathcal{C}}}{t}||\bm{\theta}||-\frac{2R_{\varphi}}{t}\right),$

### v2

@@ B.4 Proposal distribution design for the DVRPTW
The second correction factor needed is $\frac{|Q(\bm{y})|}{|Q(\bm{y}^{\prime})|}$ (see Algorithm˜2). We compute it by checking if each move is allowed, i.e., if there exists at least one $i\in V^{1}_{s}(\bm{y})$ such that $V^{2}_{s}(\bm{y})[i]\setminus\{i\}\neq\emptyset$. This can be determined in $\mathcal{O}(\mathcal{R}^{\omega})$ for each move.
Appendix C Details on the DVRPTW
C.1 Overview of the challenge.
We evaluate the proposed approach on a large-scale, ML-enriched combinatorial optimization problem: the EURO Meets NeurIPS 2022 Vehicle Routing Competition [Kool et al., 2023]. In this dynamic vehicle routing problem with time windows (DVRPTW), requests arrive continuously throughout a planning horizon, which is partitioned into a series of delivery waves $\mathcal{W}=\{\left[\tau_{0},\tau_{1}\right],\left[\tau_{1},\tau_{2}\right],\dots,\left[\tau_{|\mathcal{W}|-1},\tau_{|\mathcal{W}|}\right]\}$.
$\min_{f}\mathbb{E}\left[c_{\mathcal{W}}(f)\right],\quad\text{with}\quad c_{\mathcal{W}}(f)\coloneqq\sum_{\omega\in\mathcal{W}}c(f(\bm{x}^{\omega})),$
C.2 Reduction to supervised learning.
We follow the method of [Baty et al., 2023], which was the winning approach for the challenge. Central to this approach is the concept of prize-collecting dynamic vehicle routing problem with time windows (PC-VRPTW). In this setting, each request $i\in\mathcal{R}^{\omega}$ is assigned an artificial prize $\theta_{i}^{\omega}\in\mathbb{R}$, that reflects the benefit of serving it. The prize of the depot $D$ is set to $\theta^{\omega}_{D}=0$. The objective is then to identify a set of routes that maximizes the total prize collected while minimizing the associated travel costs. The model $g_{W}$ predicts the prize vector $\bm{\theta}^{\omega}=g_{W}(\bm{x}^{\omega})$. Denoting $\varphi(\bm{y})\coloneqq-\langle\bm{c},\bm{y}\rangle$, the corresponding optimization problem can be written as
The overall pipeline is summarized in Fig.˜1. Following [Baty et al., 2023], we approximately solve the problem in Eq.˜11 using the prize-collecting HGS heuristic (PC-HGS), a variant of hybrid genetic search (HGS) [Vidal, 2022]. We denote this approximate solver $\widetilde{\bm{y}}\approx\widehat{\bm{y}}$, so that their proposed policy decomposes as $f_{W}\coloneqq\widetilde{{\bm{y}}}\circ g_{W}$. The ground-truth routes are created by using an anticipative strategy, i.e., by solving multiple instances where all future information is revealed from the start, and the requests’ arrival times information is translated into time windows (thus removing the dynamic aspect of the problem). This anticipative policy, which we note $f^{\star}$ (which cannot be attained as it needs unavailable information) is thus the target policy imitated by the model – see Section˜C.7 for more details.
C.3 Perturbation-based baseline.
In [Baty et al., 2023], a perturbation-based method [Berthet et al., 2020] was used. This method is based on injecting noise in the PC-HGS solver $\widetilde{\bm{y}}$. Similarly to our approach, the parameters $W$ can then be learned using a Fenchel-Young loss, since the loss is minimized when the perturbed solver correctly predicts the ground truth. However, since $\widetilde{\bm{y}}$ is not an exact solver, all theoretical learning guarantees associated with this method (e.g., correctness of the gradients) no longer hold.
C.4 Proposed approach.
Our proposed approach instead uses the Fenchel-Young loss associated with the proposed layer, which is minimized when the proposed layer correctly predicts the ground-truth. At inference time, however, we use $f_{W}\coloneqq\widetilde{\bm{y}}\circ g_{W}$. We use a mixture of proposals, as defined in Algorithm˜2. To design each proposal $q_{s}$, we build randomized versions of moves specifically designed for the prize-collecting dynamic vehicle routing problem with time windows. More precisely, we base our proposals on moves used in the local search part of the PC-HGS algorithm, which are summarized in Table˜2. The details of turning these moves into proposal distributions with tractable individual correction ratios are given in Section˜B.4.
C.5 Performance metric.
C.6 Results.
In Fig.˜2, we observe that the initialization method plays an important role, and the ground-truth-based ones greatly outperform the random one.
We observe that the number of Markov iterations $K$ is an important performance factor. Interestingly, the ground-truth initialization significantly improves the learning process for small $K$.
In Table˜3, we compare training methods with fixed compute time budget for the layer (perturbed solver or proposed MCMC approach), which is by far the main computational bottleneck. This parameter limits the time allowed for a single forward pass through the combinatorial optimization layer (be it the perturbed inexact oracle or the proposed method). In both cases, the backward pass through the layer is immediate, as a property of the expression of the gradient of Fenchel-Young losses. The models are selected using a validation set and evaluated on the test set. We observe that the proposed approach significantly outperforms the perturbation-based method [Berthet et al., 2020] using $\widetilde{{\bm{y}}}$ in low time limit regimes, thus allowing for faster and more efficient training.
@@ Appendix C Details on the DVRPTW
Full experimental details and additional results on the impact of temperature are given in Section˜C.7.
C.7 Additional experimental details and results for Section˜5.1
@@
Following Baty et al. [2023], the differentiable ML model $g_{W}$ is implemented as a sparse graph neural network. We also use the same feature set, which represents the system state ${\bm{x}}^{\omega}$ as a vector comprising request-level features, such as coordinates, time windows, demands, travel time to the depot, and quantiles from the distribution of the travel time to all other requests (named complete feature set, and described in the Table 4 of their paper). We use the same training, validation, and testing datasets, which are created from $30$, $15$ and $25$ problem instances respectively. The training set uses a sample size of $50$ requests per wave, while the rest use $100$. The solutions in the training dataset, i.e., the examples from the anticipative strategy $f^{\star}$ imitated by the model, are obtained by solving the corresponding offline VRPTWs using HGS [Vidal, 2022] with a time limit of $3600$ seconds. During evaluation, the PC-HGS solver $\tilde{{\bm{y}}}$ is used with a constant time limit of $60$ seconds for all models. We use Adam [Kingma and Ba, 2017] together with the proposed stochastic gradient estimators, with a learning rate of $5\cdot 10^{-3}$. Each training is performed using only a single CPU worker. For Fig.˜2, we use a temperature $t=10^{2}$. For Table˜3, we use $1$ Monte-Carlo sample for the perturbation-based method and $1$ Markov chain for the proposed approach (in order to have a fair comparison: an equal number of oracle calls / equal compute).
@@
Each training is performed $50$ times with the same parameters and different random seeds. Then, the learning curves are averaged, and plotted with a $95\%$ confidence interval. For the results in Table˜3, we report the performance of the best model iteration (selected with respect to the validation set) on the test set. This procedure is also averaged over $50$ trainings, and reported with $95\%$ confidence intervals.
@@
In Fig.˜14, we report model performance for varying temperature $t$. Interestingly, lower temperatures perform better when using random initialization. In the ground-truth initialization setting, a sweet spot is found at $t=10^{2}$, but lower temperatures do not particularly decrease performance.
@@
Appendix D Proofs
D.1 Proof of Eq.˜4
@@ Appendix D Proofs
At fixed temperature $t_{k}=t$, the iterates of Algorithm˜1 (MH case) follow a time-homogenous Markov chain, defined by the following transition kernel $P_{\bm{\theta},t}$:
@@
$P_{\bm{\theta},t}(\bm{y},\bm{y})$ $=1-\hskip-10.00002pt\sum_{\bm{y}^{\prime}\in\mathcal{N}(\bm{y})}P_{\bm{\theta},t}(\bm{y},\bm{y}^{\prime})$
@@
Thus, we have $P_{\bm{\theta},t}(\bm{y},\bm{y})>0$, which implies that the chain is aperiodic. As an irreducible and aperiodic Markov Chain on a finite state space, it converges to its stationary distribution and the latter is unique [Freedman, 2017]. Finally, one can easily check that the detailed balance equation is satisfied for ${\pi}_{\bm{\theta},t}$, i.e.:
@@
$\xrightarrow[t\rightarrow\infty]{}\frac{1}{|\mathcal{Y}|},$
as $\exp(x/t)\xrightarrow[t\rightarrow\infty]{}1$ for all $x\in\mathbb{R}$. Thus, ${\pi}_{\bm{\theta},t}$ converges to the uniform distribution on $\mathcal{Y}$, and its expectation converges to the average of all structures.
Expression of the Jacobian. Let $A_{t}:\bm{\theta}\mapsto t\cdot\log\sum_{\bm{y}\in\mathcal{Y}}\exp\left(\langle\bm{\theta},\bm{y}\rangle+\varphi(\bm{y})\right)$ be the cumulant function of the exponential family defined by ${\pi}_{\bm{\theta},t}$, scaled by $t$. One can easily check that we have $\nabla_{\bm{\theta}}A_{t}(\bm{\theta})={\widehat{\bm{y}}_{t}}(\bm{\theta})$. Thus, we have $J_{\bm{\theta}}{\widehat{\bm{y}}_{t}}(\bm{\theta})=\nabla^{2}_{\bm{\theta}}A_{t}(\bm{\theta})$. However, we also have that the hessian matrix of the cumulant function $\bm{\theta}\mapsto\frac{1}{t}A_{t}(\bm{\theta})$ is equal to the covariance matrix of the random vector $\frac{Y}{t}$ under ${\pi}_{\bm{\theta},t}$ [Wainwright and Jordan, 2008]. Thus, we have:
@@
Let $K_{\bm{\theta},t}$ be the Markov transition kernel associated to Algorithm˜2, which can be written as:
$K_{\bm{\theta},t}(\bm{y},\bm{y}^{\prime})=\begin{cases}\sum_{\begin{subarray}{c}s\in Q(\bm{y})\\ \text{s.t. }q_{s}(\bm{y},\bm{y}^{\prime})>0\end{subarray}}\frac{1}{|Q(\bm{y})|}q_{s}(\bm{y},\bm{y}^{\prime})\min\left(1,\frac{|Q(\bm{y})|}{|Q(\bm{y}^{\prime})|}\cdot\frac{q_{s}(\bm{y}^{\prime},\bm{y}){\pi}_{{\bm{\theta}},t}(\bm{y}^{\prime})}{q_{s}(\bm{y},\bm{y}^{\prime}){\pi}_{{\bm{\theta}},t}(\bm{y})}\right)&\text{if $\bm{y}^{\prime}\in\bar{\mathcal{N}}(\bm{y})$,}\\ 1-\sum_{\bm{y}^{\prime\prime}\in\bar{\mathcal{N}}(\bm{y})}K_{\bm{\theta},t}(\bm{y},\bm{y}^{\prime\prime})&\text{if $\bm{y}^{\prime}=\bm{y}$,}\\ 0&\text{else.}\end{cases}$
@@
Thus, we only have to check that the detailed balance equation
$\pi_{{\bm{\theta}},t}(\bm{y})K_{\bm{\theta},t}(\bm{y},\bm{y}^{\prime})={\pi}_{{\bm{\theta}},t}(\bm{y}^{\prime})K_{\bm{\theta},t}(\bm{y}^{\prime},\bm{y})$
is satisfied for all $\bm{y}^{\prime}\in\bar{\mathcal{N}}(\bm{y})$. We have:
${\pi}_{{\bm{\theta}},t}(\bm{y})K_{\bm{\theta},t}(\bm{y},\bm{y}^{\prime})=\sum_{\begin{subarray}{c}s\in Q(\bm{y})\\ \text{s.t. }q_{s}(\bm{y},\bm{y}^{\prime})>0\end{subarray}}\biggl[\frac{q_{s}(\bm{y},\bm{y}^{\prime}){\pi}_{{\bm{\theta}},t}(\bm{y})}{|Q(\bm{y})|}\min\left(1,\frac{|Q(\bm{y})|}{|Q(\bm{y}^{\prime})|}\cdot\frac{q_{s}(\bm{y}^{\prime},\bm{y}){\pi}_{{\bm{\theta}},t}(\bm{y}^{\prime})}{q_{s}(\bm{y},\bm{y}^{\prime}){\pi}_{{\bm{\theta}},t}(\bm{y})}\right)\biggr].$
The main point consists in noticing that the undirectedness assumption for each individual neighborhood graph $G_{\mathcal{N}_{s}}$ implies:
@@
Thus, a simple case analysis on how $|Q(\bm{y})|q_{s}(\bm{y}^{\prime},\bm{y}){\pi}_{{\bm{\theta}},t}(\bm{y}^{\prime})$ and $|Q(\bm{y}^{\prime})|q_{s}(\bm{y},\bm{y}^{\prime}){\pi}_{{\bm{\theta}},t}(\bm{y})$ compare allows us to observe that the expression of ${\pi}_{{\bm{\theta}},t}(\bm{y})K_{\bm{\theta},t}(\bm{y},\bm{y}^{\prime})$ is symmetric in $\bm{y}$ and $\bm{y}^{\prime}$, which concludes the proof. ∎
D.4 Proof of strict convexity
@@
[…]
@@
The proposed Fenchel-Young loss can also be obtained via distribution-space regularization. Let $s_{\bm{\theta}}\coloneqq\left(\langle\bm{\theta}\,,\,\bm{y}\rangle+\varphi(\bm{y})\right)_{\bm{y}\in\mathcal{Y}}\in\mathbb{R}^{|\mathcal{Y}|}$ be a vector containing the score of all structures, and $L_{-tH}:\mathbb{R}^{|\mathcal{Y}|}\times\Delta^{|\mathcal{Y}|}\rightarrow\mathbb{R}$ be the Fenchel-Young loss generated by tH-tH, where $H$ is the Shannon entropy. We have $\nabla_{s_{\bm{\theta}}}(-tH)^{*}(s_{\bm{\theta}})={\pi}_{\bm{\theta},t}$. The chain rule further gives $\nabla_{\bm{\theta}}(-tH)^{*}(s_{\bm{\theta}})=\mathbb{E}_{{\pi}_{\bm{\theta},t}}\left[Y\right]$. Thus, we have $\nabla_{\bm{\theta}}L_{-tH}(s_{\bm{\theta}}\,;\bm{p}_{\bm{y}})=\nabla_{\bm{\theta}}\ell_{t}(\bm{\theta}\,;\bm{y})$, where $\bm{p}_{\bm{y}}$ is the Dirac distribution on $\bm{y}$. In the case where $\varphi\equiv 0$ and $t=1$, we have $\Omega_{t}(\bm{\mu})=-\left(\max_{\bm{p}\in\Delta^{|\mathcal{Y}|}}H^{s}(\bm{p})\text{ s.t. }\mathbb{E}_{\bm{p}}\left[Y\right]=\bm{\mu}\right)$, with $H^{s}$ the Shannon entropy [Blondel et al., 2020], and $\ell_{t}$ is known as the CRF loss [Lafferty et al., 2001].
D.5 Proof of Section˜4.4
@@
Moreover, we already derived $\nabla^{2}_{\bm{\theta}}A_{t}(\bm{\theta}_{0})=\frac{1}{t}\operatorname{\mathsf{cov}}_{{\pi}_{\bm{\theta}_{0},t}}\left[Y\right]$ in Section˜D.2, leading to the simplified asymptotic normality given in the proposition.
D.6 Proof of Section˜4.4
@@
[…]
A known result gives $\lambda_{\bm{\theta}}\leq\rho(\bm{\theta})$ with $\rho(\bm{\theta})=\max_{\lambda\in S_{\bm{\theta}}\backslash\{1\}}|\lambda|$ [Madras and Randall, 2002], where $S_{\bm{\theta}}$ is the spectrum of the transition kernel $P_{\bm{\theta},t}$ (in this context, $1-\rho(\bm{\theta})$ is known as the spectral gap of the Markov chain). To bound $\rho(\bm{\theta})$, we use the results of Ingrassia [1994], which study the Markov chain with transition kernel $P^{\prime}_{\bm{\theta},t}$, such that $P_{\bm{\theta},t}=\frac{1}{2}\left(I+P^{\prime}_{\bm{\theta},t}\right)$. It corresponds to the same algorithm, but with a proposal distribution $q^{\prime}$ defined as:
@@
As $P^{\prime}_{\bm{\theta},t}$ is a row-stochastic matrix, Gershgorin’s circle theorem gives that its spectrum is included in the complex unit disc. Moreover, one can easily check that the associated Markov chain is also reversible with respect to ${\pi}_{\bm{\theta},t}$, and the corresponding detailed balance equation gives:
$\forall\bm{y},\bm{y}^{\prime}\in\mathcal{Y},\;{\pi}_{\bm{\theta},t}(\bm{y})P^{\prime}_{\bm{\theta},t}(\bm{y},\bm{y}^{\prime})={\pi}_{\bm{\theta},t}(\bm{y}^{\prime})P^{\prime}_{\bm{\theta},t}(\bm{y}^{\prime},\bm{y}),$
@@
$\forall\bm{y},\bm{y}^{\prime}\in\mathcal{Y},\;\sqrt{\frac{{\pi}_{\bm{\theta},t}(\bm{y})}{{\pi}_{\bm{\theta},t}(\bm{y}^{\prime})}}P^{\prime}_{\bm{\theta},t}(\bm{y},\bm{y}^{\prime})=\sqrt{\frac{{\pi}_{\bm{\theta},t}(\bm{y}^{\prime})}{{\pi}_{\bm{\theta},t}(\bm{y})}}P^{\prime}_{\bm{\theta},t}(\bm{y}^{\prime},\bm{y})$
@@
$\Pi_{\bm{\theta}}^{1/2}P^{\prime}_{\bm{\theta},t}\Pi_{\bm{\theta}}^{-1/2}=\Pi_{\bm{\theta}}^{-1/2}P^{\prime\top}_{\bm{\theta},t}\Pi_{\bm{\theta}}^{1/2},$
where $\Pi_{\bm{\theta}}=\operatorname{\mathsf{diag}}({\pi}_{\bm{\theta};t})$. Thus, the matrix $\Pi_{\bm{\theta}}^{1/2}P^{\prime}_{\bm{\theta},t}\Pi_{\bm{\theta}}^{-1/2}$ is symmetric, and the spectral theorem ensures its eigenvalues are real. As it is similar to the transition kernel $P^{\prime}_{\bm{\theta},t}$ (with change of basis matrix $\Pi_{\bm{\theta}}^{-1/2}$), they share the same spectrum $S^{\prime}_{\bm{\theta}}$, and we have $S^{\prime}_{\bm{\theta}}\subset\left[-1,1\right]$. Let us order $S^{\prime}_{\bm{\theta}}$ as $1\leq\lambda^{\prime}_{\min}\leq\dots\leq\lambda^{\prime}_{2}\leq\lambda^{\prime}_{1}=1$. As $P_{\bm{\theta},t}=\frac{1}{2}\left(I+P^{\prime}_{\bm{\theta},t}\right)$, we clearly have $\rho(\bm{\theta})=\frac{1+\lambda^{\prime}_{2}}{2}$. Thus, we can use Theorem 4.1 of Ingrassia [1994], which gives $\lambda^{\prime}_{2}\leq 1-G\cdot Z(\bm{\theta})\exp(-m\left(\bm{\theta})\right)$ (we keep their notations for $Z$ and $m$, and add the dependency in $\bm{\theta}$ for clarity), where $G$ is a constant depending only on the graph $G_{\mathcal{N}}$, and with:
@@
[…]